2.It is 2018 and Mark has the following investments. He owns 40% of Mav & Associates. Mav is an accounting firm and distributes to Mark $50,000. The firm has sales income of 1.600,000 and operating expenses of $1,100,000. The firm has property with an adjusted basis of $800,000 and paid W2 wages for the year of $500,000 Marks is married and his taxable income lexcluding any income from Mav &Associates is $450,000. How will Marks personal tax return be affected if? (5.5 points) a. Mav is a C-Corporation b. Mav is a Partnership c. Mav is an LLC, if you need more information please indicate what information you would need. d. How would your answer change in part b if Mav & Assaciates was an engineering firm instead of an accounting firm?
Answer #1

Answer for a)

sales= 1600000

   expences= 1100000

   profit =1600000-1100000=500000

CORPORATION TAX @20% =500000*20/100=100000

Share of Mark = 100000*40/100=40000

Therefore Mark's taxable income is 40000+450000=490000

Marks personal tax is

Upto 2.5 lakh    NILL

above 2.5 240000*10/100=24000

Mark's personal tax return =24000

Answer for b)

   sales= 1600000

   expences= 1100000

   book profit =1600000-1100000=500000

BOOK PROFIT AMOUNT DEDUCTIBLE u/s 40B

ABOVE 3 LAKH 60% OF BOOK PROFIT

Therefore taxable income is 500000*60/100= 300000

hence share of Mark is 300000*40/100=120000

Therefore Mark's taxable income is 120000+450000=570000

TAX SLAB    TAX RATE   

Upto 2.5 lakh    NILL

2.5 TO 5 LAKH    10%

ABOVE 20%

Marks personal tax is

Upto 2.5 lakh    NILL

2.5    250000*10/100=25000

70000(570000-500000) 70000*20/100=14000

Mark's personal tax return =25000+14000=39000

Answer for c)

LLC stands for Limited liability company. It is not a taxing entity and it is not recognized by IRS . The IRS says LLC may be taxed as a partnership firm. Hence calculation of tax return is same as partnership firm.

Answer for d )

Since taxable income is less than 1000000 , Marks tax return is not affected.